OpenStudy (anonymous):

Suppose that a substance decays from 10g to 3g in 20 days. Show how to find the half-life of the substance.

OpenStudy (anonymous):

3=10(.5)^t solve for t 3/10=.5^t log(3/10)=t*log(.5) solve for t. Hope you can do that.

OpenStudy (anonymous):

I came up with 1.7372 does that look right to you?

OpenStudy (anonymous):

yeah. good job.